OpenStudy (anonymous):

Find the quotient of 21x^2y^6+6xy^3-30xy over 3xy A. 21x^2y^6+6xy^3-10 B. 7xy^5+2y^2+10 C. 7xy^5+2y^2-10 D. 7xy^5+2y-10

OpenStudy (anonymous):

\[\frac{ x^n }{ x }=x^(n-1)\]

OpenStudy (anonymous):

I divided each one my the bottom number 3

OpenStudy (anonymous):

by*

OpenStudy (perl):

err, factor out 3xy from the numerator

OpenStudy (anonymous):

for the first i got 7xy^5 + i kept the sign then i got 2y then 30 divded by 3 is 10 so i got the answer which is D

OpenStudy (perl):

nope, i dont think its D

OpenStudy (perl):

i think its C

OpenStudy (anonymous):

How?...

OpenStudy (anonymous):

6xy^3 over by 3xy^0

OpenStudy (anonymous):

the 0 is 1 0fcourse so 3-1 is 2

OpenStudy (anonymous):

that was my mistake i forgot the exponent

OpenStudy (perl):

21x^2y^6+6xy^3-30xy over 3xy